# Time, Travel, and Energy in the Uniform Dispersion Problem

### v1

@@ V-A Analysis
The execution of (synchronous) FCDFS over a region of interest $R$ is deterministic, and so we know in advance the path each robot shall take. Let us define a graph, $G(R)$, whose nodes are the locations of $R$ and where there is an edge between two locations $v_{i},v_{j}$ if at some point a robot moves from $v_{i}$ to $v_{j}$ during an execution of FCDFS. Recalling our analysis of FCDFS, it is not difficult to prove that $G(R)$ is a tree. Although robot wake-up times are random in our asynchronous model, the ultimate path each robot follows according to AsynchFCDFS is deterministic - it is the same path that robot would follow under FCDFS. Hence, in AsynchFCDFS, too, each robot moves down the edges of the tree $G(R)$.
@@
Assume first, for contradiction, that (a) is true and (b) is not true. By the structure of $P(n)$, $A_{i}^{P(n)}$ being settled means that $depth_{i}^{P(n)}(t)=n-i-1$ and that (if $i>1$) all of $v_{n-i},\ldots v_{n-1}$ contain the settled robots $A_{1}^{P(n)},\ldots A_{i-1}^{P(n)}$ at time $t$. By (1) and (2), this means that $A_{1}^{R},\ldots A_{i-1}^{R}$ are settled at time $t$. Since $A_{i}^{R}$ is not settled at time $t+1$, this must mean that either it moved and increased its depth since time $t$, or it was adjacent a active agent in one of its primary directions at time $t$. The latter is impossible, since (by the FCDFS algorithm) such an agent must have been one of the agents $A_{1}^{R},\ldots A_{i-1}^{R}$. Hence, $A_{i}^{R}$ must have moved at time $t$. By our assumptions, this implies $depth_{i}^{R}(t+1)\geq depth_{i}^{P(n)}(t)+1=n-i$. But this is impossible, since there are at least $i-1$ settled agents in $R$ at time $t+1$, hence the maximum depth an active agent can have is $n-i-1$. Contradiction.
@@ VI Empirical Analysis
To verify our theoretical findings, we ran numerical simulations on simply-connected environments of different sizes and obstacle placements. To measure the improvement attained by FCDFS and AsynchFCDFS, we compared it to Hsiang et al.’s DFLF and BFLF algorithms [3]. DFLF is described in Proposition III.4. BFLF (Breadth-First Leader-Follower) is a different makespan-optimal dispersion algorithm that seeks to spread out the agents in a more even way by having multiple leaders, reducing total and maximal individual travel at the cost of energy. Table II summarizes simulation results in two different environments. We see that (as anticipated) the travel for FCDFS and AsynchFCDFS is identical, but makespan and energy use are larger for smaller $p$. Additionally, AsynchFCDFS’s average performance is below (i.e., better than) the bounds of Propositions V.1 and V.2, and V.3 for the tested $p$ values.
@@
Next, we were interested in how FCDFS and AsynchFCDFS’s performance varies as $n$ (the environment size) grows. We plotted the performance of these algorithms alongside DFLF and BFLF on square grids of increasing size: see Figure 9. We see that, although the synchronous algorithms DFLF, BFLF and FCDFS are matched in terms of makespan, FCDFS attains significantly better energy and travel. The asynchronous variant of FCDFS attains worse makespan, as expected (since robots are inactive in some timesteps), but still attains better energy and travel than DFLF and BFLF for $p=0.5,0.75$. AsynchFCDFS’s maximal individual energy use and makespan are below the bounds anticipated by Propositions V.1 and V.2, and total energy use is below the bound anticipated by V.3.
TABLE II: Simulation results over the environments shown in Figures 4 and 5. Each measured performance metric written in the Table is the rounded average of $10$ trials.
Experiment Name $\boldsymbol{T_{\textbf{total}}}$ $\boldsymbol{T_{\textbf{max}}}$ $\boldsymbol{E_{\textbf{total}}}$ $\boldsymbol{E_{\textbf{max}}}$ $\boldsymbol{\mathcal{M}}$
FCDFS, Fig. 4 Env. ($n=669$) 35 1033510335\,10335 103 $99$ 35 7723577235\,77235 772 $100$ $1338$
AsynchFCDFS ($p=0.75$), Fig. 4 Env. 35 1033510335\,10335 103 $99$ 69 0926909269\,09269 092 $137$ $2565$
AsynchFCDFS ($p=0.5$), Fig. 4 Env. 35 1033510335\,10335 103 $99$ 118 658118658118\,658118 658 $171$ $4350$
BFLF, Fig. 4 Env. 39 6493964939\,64939 649 $117$ 122 787122787122\,787122 787 $197$ $1338$
DFLF, Fig. 4 Env. 91 9599195991\,95991 959 $260$ 92 6289262892\,62892 628 $261$ $1338$
FCDFS, Fig. 5 Env. ($n=277$) $5909$ $38$ $6186$ $39$ $554$
AsynchFCDFS ($p=0.75$), Fig. 5 Env. $5909$ $38$ 11 7831178311\,78311 783 $46$ $1027$
AsynchFCDFS ($p=0.5$), Fig. 5 Env. $5909$ $38$ 20 0012000120\,00120 001 $68$ $1725$
BFLF, Fig. 5 Env. $7313$ $56$ 28 3582835828\,35828 358 $130$ $554$
DFLF, Fig. 5 Env. 17 2141721417\,21417 214 $134$ 17 4911749117\,49117 491 $135$ $554$
@@
Figure 9: Comparative results of the algorithms across different metrics. For each $k\in[1,30]$, we generated $40$ different square grid environments of size $n=k\times k$ with randomly placed source locations. Each data point is the average of $40$ simulations of its respective algorithm, one per grid environment. In each graph, the $x$ axis is $n$ and the $y$ axis is the performance measure of interest. Note that some algorithms overlap perfectly on metrics such as makespan.
@@ VII Conclusion
Our results shed light on the relationships between the capabilities of robots, their environment, and the quality of the collective behavior they can achieve. The fact that robots can minimize the travel or makespan even with very limited capabilities, but cannot minimize energy even with arbitrarily strong capabilities, attests to the difficulty of designing energy-efficient collective behaviors. On the other hand, the fact that energy can be minimized by ant-like robots in common types of environments suggests energy efficiency in swarm robotic systems may be practically attainable, especially in restricted settings. Broadly, our results point to the importance of co-design approaches in swarm robotics, carefully tailoring the capabilities of robots to the specific features of the environment they are expected to operate in.
@@ Acknowledgements
The authors are very grateful to Ofer Zeitouni (Weizmann Institute of Science) for helpful discussions.
@@
Informally, Proposition III.9 shows that decentralized algorithms with sufficient memory and communication capabilities are able to simulate a fully centralized algorithm at the cost of a multiplicative ($\Delta$) time delay, as long as the robots’ communication network remains connected. In other words, as long as the network of decentralized robots collectively senses the same environmental data the target centralized algorithm does, its decision-making capabilities are just as strong. We claim that ant-like algorithms (algorithms requiring small, constant visibility and persistent state memory, and no communication, to be run on their target class of environments) are “strongly” decentralized: they cannot make the same type of decisions a centralized algorithm can, even when collectively they sense the same environmental data. We prove this by giving an example of a centralized, visibility preserving algorithm that ant-like algorithms cannot simulate, even assuming arbitrary $\Delta$-delay.
@@
-B FCDFS Using 5 Bits of Memory
Algorithm 3 5-bit FCDFS
Let $v$ be the current location of $A$.
if all neighboring vertices of $v$ are occupied then
     Settle.
else if $b_{4}b_{5}=00$ then
     $b_{4}b_{5}\leftarrow 10$
end if
if $A$ cannot move in primary or secondary directions then
     if $v$ has just one unoccupied neighbour then
         Settle.
     else if $(b_{5}=1\land b_{3}+b_{4}=1)$lor $diag(v)$ is unoccupied then
         Settle.
     else
         Set primary direction to obstacle-less direction not equal to $180$-degree rotation of previous direction stepped in (i.e. the neighbour of $v$ we haven’t visited yet; this can be inferred from $b_{1}b_{2}$ and $b_{3}$).
         $b_{4}b_{5}\leftarrow 10$
     end if
end if
if $b_{4}b_{5}$ was not updated at this time step $\triangleright$ i.e. $b_{5}=1$ or time to update $b_{5}$
     $b_{4}b_{5}\leftarrow b_{3}1$
end if
if $A$ can move in its primary direction then
     Step in the primary direction.
     $b_{3}\leftarrow 0$
else if $A$ can step in secondary direction then
     Step in the secondary direction.
     $b_{3}\leftarrow 1$
else
     Settle.
end if
FCDFS can be implemented using 5 bit of memory - see Algorithm 3. In this implementation, a robot’s state is described by bits $b_{1}b_{2}b_{3}b_{4}b_{5}$. All bits are initially 00. $b_{1}b_{2}$ describe the primary direction (one of four), and $b_{3}$ tells us whether the previous step was taken in the primary direction (if $b_{3}=0$) or in the secondary direction (if $b_{3}=1$). $b_{4}b_{5}$ is a counter that is reset to $10$ upon entering a hall or one step after initialization, and thereafter is equal to $*1$, where * is a bit that tells us whether we walked in the primary or secondary direction two steps ago (by copying $b_{3}$). A robot that detects an obstacle at its diagonal interprets its position as a fake hall (i.e. a corner) as long as $b_{5}=1$ and $b_{3}+b_{4}=1$, that is, as long as at least one step passed since the last hall, and our position two steps ago was diagonal to us.

### v2

@@ V-A Analysis
The execution of (synchronous) FCDFS over a region of interest $R$ is deterministic, and so we know in advance the path each robot will take. Let us define a graph, $G(R)$, whose nodes are the locations of $R$ and where there is an edge between two locations $v_{i},v_{j}$ if at some point a robot moves from $v_{i}$ to $v_{j}$ during an execution of FCDFS. Recalling our analysis of FCDFS, it is not difficult to prove that $G(R)$ is a tree. Although robot wake-up times are random in our asynchronous model, the ultimate path each robot follows according to AsynchFCDFS is deterministic–the same path it would follow under FCDFS. Hence, in AsynchFCDFS, too, each robot moves down the edges of the tree $G(R)$.
@@
Assume first, for contradiction, that (a) is true and (b) is not true. By the structure of $P(n)$, $A_{i}^{P(n)}$ being settled means that $depth_{i}^{P(n)}(t)=n-i-1$ and that (if $i>1$) all of $v_{n-i},\ldots v_{n-1}$ contain the settled robots $A_{1}^{P(n)},\ldots A_{i-1}^{P(n)}$ at time $t$. By (1) and (2), this means that $A_{1}^{R},\ldots A_{i-1}^{R}$ are settled at time $t$. Since $A_{i}^{R}$ is not settled at time $t+1$, this must mean that either it moved and increased its depth since time $t$, or it was adjacent an active agent in one of its primary directions at time $t$. The latter is impossible, since (by the FCDFS algorithm) such an agent must have been one of the agents $A_{1}^{R},\ldots A_{i-1}^{R}$. Hence, $A_{i}^{R}$ must have moved at time $t$. By our assumptions, this implies $depth_{i}^{R}(t+1)\geq depth_{i}^{P(n)}(t)+1=n-i$. But this is impossible, since there are at least $i-1$ settled agents in $R$ at time $t+1$, hence the maximum depth an active agent can have is $n-i-1$. Contradiction.
@@ VI Empirical Analysis
To verify our theoretical findings and measure performance metrics, we simulated our algorithms on various simply-connected environments. The simulator implements our model (Section II), and implements FCDFS and AsynchFCDFS using their minimal required capabilities ($(2,0,5)$ and $(2,1,5)$ respectively). To measure the improvement attained by FCDFS and AsynchFCDFS, we compared it to Hsiang et al.’s DFLF and BFLF algorithms [3]. DFLF is described in Proposition III.4. BFLF (Breadth-First Leader-Follower) is a different makespan-optimal dispersion algorithm that seeks to spread out the agents more evenly by having multiple leaders, reducing total and maximal individual travel at the cost of energy.
We tested our algorithms on several environment types: custom designed environments (Figures 4 and 5), realistic environments from the ”Benchmarks for Grid-Based Pathfinding” repository [50], and systematically generated square grids of increasing size. Table II summarizes simulation results across these environments. The benchmark environments, shown in Figure 10, include complex layouts and indoor urban environments that we selected and minimally modified to ensure simple-connectedness. This diverse set of environments allows us to validate our theoretical results under varying practical conditions.
Figure 10: Benchmark environments based on [50].
As guaranteed by our theoretical results, FCDFS achieves optimal performance across all tested environments, including the realistic benchmarks. The performance gap between FCDFS and previous algorithms (DFLF, BFLF) remains consistent, and large, across environment types, with no cases where previous methods outperform FCDFS in any of our metrics. However, while FCDFS achieves optimality in simply connected environments, DFLF and BFLF remain valuable options for environments with holes or obstacles that break simple connectedness.
Looking at specific performance patterns, we observe that the travel distance for FCDFS and AsynchFCDFS is identical, but makespan and energy use are larger for smaller $p$. For $p=0.5$ and $p=0.75$, AsynchFCDFS’s average performance is consistent with the theoretical bounds of Propositions V.1 and V.2, and V.3. While our theoretical analysis holds, provably, for all $p\in(0,1]$, these results provide further empirical validation of our claims.
@@
Next, we were interested in how FCDFS and AsynchFCDFS’s performance varies as $n$ (the environment size) grows. We plotted the performance of these algorithms alongside DFLF and BFLF on square grids of increasing size: see Figure 11. We see that, although the synchronous algorithms DFLF, BFLF and FCDFS are matched in terms of makespan, FCDFS attains significantly better energy and travel. AsynchFCDFS attains worse makespan, as expected (since robots are inactive in some timesteps), but still attains better energy and travel than DFLF and BFLF for $p=0.5,0.75$. AsynchFCDFS’s maximal individual energy use and makespan are consistent with the theoretical bounds given by Propositions V.1 and V.2, and total energy use consistent with V.3.
TABLE II: Simulation results across various environments. Metrics represent averages over 10 trials. Abbreviations: ’k’ indicates thousands, and ’AsFCDFS’ refers to Asynchronous FCDFS. Standard deviations are indicated by ±. If no ± is shown, the standard deviation is 0.
[…]
Figure 10(a) ($n=2501$) FCDFS 78k 61 80k 62 5002
AsFCDFS (p=0.75) 78k 61 153k ±$\pm 2k$ 94 ±$\pm 17$ 9480 ±$\pm 47$
AsFCDFS (p=0.5) 78k 61 261k ±$\pm 4k$ 97 ±$\pm 33$ 16034 ±$\pm 91$
BFLF 90k ±$\pm 2k$ 82 ±$\pm 7$ 1406k ±$\pm 100k$ 269 ±$\pm 470$ 5002
DFLF 895k ±$\pm 127k$ 647 ±$\pm 110$ 897k ±$\pm 127k$ 648 ±$\pm 110$ 5002
Figure 10(b) ($n=1846$) FCDFS 70k 69 72k 70 3692
AsFCDFS (p=0.75) 70k 69 140k ±$\pm 2k$ 88 ±$\pm 12$ 7073 ±$\pm 41$
AsFCDFS (p=0.5) 70k 69 236k ±$\pm 3k$ 122 ±$\pm 41$ 11922 ±$\pm 64$
BFLF 89k ±$\pm 5k$ 94 ±$\pm 9$ 691k ±$\pm 40k$ 276 ±$\pm 239$ 3692
DFLF 457k ±$\pm 81k$ 479 ±$\pm 59$ 459k ±$\pm 81k$ 480 ±$\pm 59$ 3692
Figure 10(c) ($n=1006$) FCDFS 67k 124 68k 125 2012
AsFCDFS (p=0.75) 67k 124 131k ±$\pm 2k$ 186 ±$\pm 27$ 3899 ±$\pm 17$
AsFCDFS (p=0.5) 67k 124 223k ±$\pm 3k$ 241 ±$\pm 76$ 6603 ±$\pm 56$
BFLF 72k ±$\pm 2k$ 135 ±$\pm 3$ 312k ±$\pm 12k$ 200 ±$\pm 50$ 2012
DFLF 138k ±$\pm 6k$ 252 ±$\pm 15$ 139k ±$\pm 6k$ 253 ±$\pm 15$ 2012
Figure 10(d) ($n=2122$) FCDFS 49k 45 51k 46 4244
AsFCDFS (p=0.75) 49k 45 97k ±$\pm 872$ 60 ±$\pm 9$ 7940 ±$\pm 49$
AsFCDFS (p=0.5) 49k 45 164k ±$\pm 2k$ 77 ±$\pm 30$ 13428 ±$\pm 75$
BFLF 61k ±$\pm 3k$ 74 ±$\pm 12$ 1525k ±$\pm 68k$ 417 ±$\pm 660$ 4244
DFLF 1186k ±$\pm 98k$ 961 ±$\pm 78$ 1188k ±$\pm 98k$ 962 ±$\pm 78$ 4244
Figure 10(e) ($n=6862$) FCDFS 579k 177 586k 178 13724
AsFCDFS (p=0.75) 579k 177 1143k ±$\pm 7k$ 250 ±$\pm 45$ 26782 ±$\pm 57$
AsFCDFS (p=0.5) 579k 177 1941k ±$\pm 14k$ 382 ±$\pm 143$ 45528 ±$\pm 115$
BFLF 706k ±$\pm 22k$ 215 ±$\pm 14$ 6594k ±$\pm 395k$ 521 ±$\pm 225$ 13724
DFLF 5640k ±$\pm 665k$ 1542 ±$\pm 174$ 5647k ±$\pm 665k$ 1543 ±$\pm 174$ 13724
Figure 10(f) ($n=4552$) FCDFS 404k 208 408k 209 9104
AsFCDFS (p=0.75) 404k 208 799k ±$\pm 4k$ 260 ±$\pm 59$ 17652 ±$\pm 53$
AsFCDFS (p=0.5) 404k 208 1363k ±$\pm 10k$ 478 ±$\pm 133$ 29958 ±$\pm 58$
BFLF 506k ±$\pm 25k$ 271 ±$\pm 8$ 2671k ±$\pm 127k$ 368 ±$\pm 65$ 9104
DFLF 2845k ±$\pm 656k$ 1204 ±$\pm 233$ 2850k ±$\pm 656k$ 1205 ±$\pm 233$ 9104
Figure 4 ($n=669$) FCDFS 35k 99 36k 100 1338
AsFCDFS (p=0.75) 35k 99 69k ±$\pm 1k$ 144 ±$\pm 35$ 2570 ±$\pm 19$
AsFCDFS (p=0.5) 35k 99 117k ±$\pm 3k$ 199 ±$\pm 75$ 4332 ±$\pm 53$
BFLF 40k ±$\pm 1k$ 117 ±$\pm 8$ 126k ±$\pm 8k$ 186 ±$\pm 61$ 1338
DFLF 98k ±$\pm 11k$ 276 ±$\pm 34$ 98k ±$\pm 11k$ 277 ±$\pm 34$ 1338
Figure 5 ($n=277$) FCDFS 6k 38 6k 39 554
AsFCDFS (p=0.75) 6k 38 12k ±$\pm 414$ 50 ±$\pm 6$ 1028 ±$\pm 25$
AsFCDFS (p=0.5) 6k 38 20k ±$\pm 826$ 71 ±$\pm 25$ 1716 ±$\pm 38$
BFLF 8k ±$\pm 364$ 51 ±$\pm 4$ 33k ±$\pm 2k$ 120 ±$\pm 34$ 554
DFLF 17k ±$\pm 892$ 129 ±$\pm 10$ 17k ±$\pm 892$ 130 ±$\pm 10$ 554
@@
Figure 11: Comparative results of the algorithms across different metrics. For each $k\in[10,50]$, we generated $100$ regular square grid environments of size $n=k\times k$ with different randomly placed source locations. Each data point is the average of $100$ simulations of its respective algorithm, one per such grid environment. In each graph, the $x$ axis is $n$ and the $y$ axis is the performance measure of interest. Shaded regions denote standard deviations. Note that some algorithms overlap perfectly on metrics such as makespan.
@@ VII Conclusion
Our results shed light on the relationships between the capabilities of robots, their environment, and the quality of the collective behavior they can achieve. The fact that robots can minimize the travel or makespan even with very limited capabilities, but cannot minimize energy even with arbitrarily strong capabilities, attests to the difficulty of designing energy-efficient collective behaviors. On the other hand, the fact that energy can be minimized by ant-like robots in simply connected environments suggests energy efficiency in swarm robotic systems may be practically attainable, especially in restricted settings. Broadly, our results point to the importance of co-design approaches in swarm robotics, carefully tailoring the capabilities of robots to the specific features of the environment they are expected to operate in.
@@
As our goal in this work was a formal mathematical analysis of the algorithmic components of time, travel, and energy use, we assumed an idealized grid setting. However, in the real world, energy is shaped by physical realities beyond our grid model–e.g. motion uncertainty, kinematic constraints, and environmental factors such as small debris, floor friction, and slopes. The problem of handling such empirical factors, and superimposing them top of our theoretical lower bounds, is an important direction for future work.
@@ Acknowledgements
The authors would like to thank Ofer Zeitouni (Weizmann Institute of Science) for helpful discussions.
@@
Proposition III.9 establishes that decentralized algorithms, given sufficient memory and communication capabilities, can emulate centralized algorithms while incurring only a multiplicative time delay ($\Delta$), provided their communication network remains connected. Thus, when decentralized robots collectively gather the same environmental data as their centralized counterpart, they can achieve equivalent decision-making capabilities. We claim that ant-like algorithms–algorithms requiring small, constant visibility and persistent state memory, and no communication, to run on their target class of environments–are “strongly decentralized”: these algorithms, despite potentially sensing the same collective environmental data, cannot replicate certain decisions made by centralized algorithms. We prove this by giving an example of a centralized, visibility preserving algorithm that ant-like algorithms cannot simulate, even assuming arbitrary $\Delta$-delay.